OpenStudy (anonymous):

The formula for finding the area of a square that has a side length, s, is A = s2. If a square has an area of 40 square units, what is the length of a side?

OpenStudy (anonymous):

A= 40 s= ? 40=s2 20=s

OpenStudy (anonymous):

its s squared

OpenStudy (anonymous):

oh, then you do square root of 40

OpenStudy (anonymous):

so its 

OpenStudy (anonymous):

It's \[A= 40\] \[s= ?\] \[40=s^2\] \[\sqrt{40} =s\]

OpenStudy (anonymous):

You're correct.

OpenStudy (anonymous):

thanks ;)